# Optimal Dispatch of Electricity and Water in Renewable-Integrated Desalination Plants

### v1

@@ II-D Profit-Maximizing WDP Dispatch
The WDP jointly dispatches thermal and RO units to maximize profit under capacity and energy balance constraints; hence, it solves
@@
where $\cal W$$\in\mathbb{R}_{+}$ is the total demand of desalinated water. We assume that the WDP is sized so that the minimum output of desalinated water is no less than that demanded by the water utility, i.e., $\underline{w}_{h}+\underline{w}_{r}\geq{\cal W}$. The dispatch problem in (II-D) is well-posed with global optima, as the objective is concave (Lemma 1) over a convex and closed set.
@@ III-A Optimal WDP Dispatch
Theorem 1 establishes that the WDP dispatches the TDP and RODP by measuring renewable power generation and comparing it against four thresholds defined in (6), as illustrated in Figure 3. The following corollary formalizes the relationship among these thresholds.
@@
Although Theorem 1 specifies only the optimal water dispatch, it also implicitly dispatches electricity consumption and production, given the RODP conversion function $f_{r}(\cdot)$ and the TDP water-to-power production ratio $\eta_{h}$.
@@ IV-A Special Technology Cases: RODP-Only and TDP-Only
In the RODP-only case, the WDP behaves as a controllable load under net metering, arbitraging import and export prices [Alahmed&Tong:22IEEETSG]. Whereas, in the TDP-only case, the WDP is always a producer of both water and power. Consequently, the optimal operating point, being price dependent, remains constant.
@@ V Numerical Results
Typically, the TDP operating cost function is assumed to be convex quadratic $C_{h}(p_{h})=ap^{2}_{h}+bp_{h}+c$, with $a=\mathdollar 0.008/\text{MBTU}^{2}$, $b=\mathdollar 2$/MBTU, and $c=\mathdollar 0$ [Santhosh&Farid&Toumi:14AE]. The TDP and RODP flowrate limits were set as follows $\overline{w}_{r}=8,333\;m^{3}/h,\overline{w}_{h}=3,000\;m^{3}/h,\underline{w}_{h}=\underline{w}_{r}=0\;m^{3}/h$. The conversion factors for TDP were set $\alpha_{h}=4\;m^{3}$/MBTU, $\beta_{h}=0.05$ MWh/MBTU, $\eta_{h}=\alpha_{h}/\beta_{h}=80\;m^{3}$/MWh. The conversion for RODP is $\alpha_{r}=166.67\;m^{3}$/MWh.

### v2

@@ II-D Profit-Maximizing WDP Dispatch
The WDP jointly schedules thermal and RO units to maximize profit under capacity and energy balance constraints; hence, it solves
@@
where $\cal W$$\in\mathbb{R}_{+}$ is the total demand of desalinated water. We assume that the WDP is sized so that the minimum output of desalinated water is no less than that demanded by the water utility, i.e., $\underline{w}_{h}+\underline{w}_{r}\geq{\cal W}$. The scheduling problem in (II-D) is well-posed with global optima, as the objective is concave (Lemma 1) over a convex and closed set.
@@ III-A Optimal WDP Dispatch Policy
Theorem 1 establishes that the WDP schedules the TDP and RODP by measuring renewable power generation and comparing it against four thresholds defined in (6), as illustrated in Figure 3. The following corollary formalizes the relationship among these thresholds.
@@
Although Theorem 1 specifies only optimal water dispatch, it also implicitly schedules electricity consumption and production, given the RODP conversion function $f_{r}(\cdot)$ and the TDP water-to-power production ratio $\eta_{h}$.
@@ IV-A Special Technology Cases: RODP-Only and TDP-Only
In the RODP-only case, the WDP behaves as a controllable load under net metering, arbitraging import and export prices [3]. Whereas, in the TDP-only case, the WDP is always a producer of both water and power. Consequently, the optimal operating point, being price dependent, remains constant.
@@ V Numerical Results
Typically, the TDP operating cost function is assumed to be convex quadratic $C_{h}(p_{h})=ap^{2}_{h}+bp_{h}+c$, with $a=\mathdollar 0.008/\text{MBTU}^{2}$, $b=\mathdollar 2$/MBTU, and $c=\mathdollar 0$ [26]. The TDP and RODP flowrate limits were set as follows $\overline{w}_{r}=8,333\;m^{3}/h,\overline{w}_{h}=3,000\;m^{3}/h,\underline{w}_{h}=\underline{w}_{r}=0\;m^{3}/h$. The conversion factors for TDP were set $\alpha_{h}=4\;m^{3}$/MBTU, $\beta_{h}=0.05$ MWh/MBTU, $\eta_{h}=\alpha_{h}/\beta_{h}=80\;m^{3}$/MWh. The conversion for RODP is $\alpha_{r}=166.67\;m^{3}$/MWh.